OpenStudy (anonymous):

A 50V independent voltage source supplies power to three resistors in the circuit shown...

OpenStudy (anonymous):

http://postimage.org/image/73bn26tuz/ For each resistor, Find the voltage drop and current R1= 55 Ohm R2= 380 Ohm R3= 28 Ohm

OpenStudy (anonymous):

There are two approaches we can take. 1) Find the equivalent resistance, find the voltage and current, then decompose the circuit back to it's original state. 2) Using Kirchoff's Laws Which do you prefer?

OpenStudy (anonymous):

kirchoff's law

OpenStudy (anonymous):

Good choice. Let A be the node where R3 and R1 meet; loop 1 be the loop containing the voltage source, R1 and R3; and loop 2 be the loop containing the voltage source, R2 and R3. Can you picture these without a drawing?

OpenStudy (anonymous):

yea but my mental image is a little blurry lol

OpenStudy (anonymous):

I've defined a third loop in-case we need it. Current into node A must be zero. \[0 = i_1 + i_2 - i_3\] Voltage around all loops must be zero. \[KVL_1: ~~-V_s + V_1 + V_3 = 0\]\[KVL_2:~~ -V_s + V_2 + V_3 = 0\]\[KVL_3:~~ -V_1 + V_2 = 0\]